+ What are the hypercritical points for f(x) = x3 3x? ? (lf any exist at all).
Answer #1

Given fra)= x² + x² 3x2 (n) = fens = f (a) = bat6 CA 2 Take el Ca)=0 ine bat6 =O=COM - 6 n+loor out gow? n-1 I tfs = a=4 is h